1. A gram of soil sample as received contains 15.00% water. After oven drying to remove the mixture, it was found to contain 16.00% K. Find the %K in the original sample as received. 


1
Expert's answer
2021-11-03T08:13:28-0400


16/100×1g = 0.16g

0.16+0.15= 0.31

= 0.68×100= 68%